# Low-Resource Dialect Adaptation of Large Language Models: A French Dialect Case-Study

### v2

@@ Abstract
Despite the widespread adoption of large language models (LLMs), their strongest capabilities remain largely confined to a small number of high-resource languages for which there is abundant training data. Recently, continual pre-training (CPT) has emerged as a means to fine-tune these models to low-resource regional dialects. In this paper, we study the use of CPT for dialect learning under tight data and compute budgets. Using low-rank adaptation (LoRA) and compute-efficient continual pre-training, we adapt three LLMs to the Québec French dialect using a very small dataset and benchmark them on the COLE suite. Our experiments demonstrate an improvement on the minority dialect benchmarks with minimal regression on the prestige language benchmarks with around 1% of model parameters updated. Analysis of the results demonstrate that gains are highly contingent on corpus composition. These findings indicate that CPT with parameter-efficient fine-tuning (PEFT) can narrow the dialect gap by providing cost-effective and sustainable language resource creation, expanding high-quality LLM access to minority linguistic communities. We release the first Québec French LLMs on HuggingFace.
@@ 1.  Introduction
Despite their success, most widely-used LLMs are trained on predominantly English datasets. When other languages are included in multilingual training datasets, they are collected from the high-resource prestige dialect of the language and have little coverage of low-resource regional dialects. As a result, they often struggle with local vocabulary, orthographic variants, idiomatic expressions, and code-switching phenomena that are commonly found in regional dialects. This issue, coined the dialect gap Kantharuban et al. (2023), limits their effectiveness for millions of users of these minority language varieties, thereby perpetuating inequities in access to artificial intelligence (AI) technologies.
One possible approach to addressing the dialect gap is to train LLMs using regional dialectal data. However, full model training is prohibitively expensive, while conventional fine-tuning approaches can be inefficient or lead to overfitting on relatively small dialectal datasets. Continual pre-training (CPT) on unlabeled data offers a practical compromise: by exposing models to large volumes of dialect-specific text, CPT enhances regional linguistic coverage without fully discarding the general knowledge encoded during initial pre-training (Gururangan et al., 2020a; Sarkar et al., 2022; Lee et al., 2019c).
In this paper, we explore the ability and limitations of CPT to adapt LLMs to a regional dialect using a small amount of unlabeled regional texts. As a case study, we consider specifically the case of the Québec regional dialect of French, also called Québécois. To make adaptation feasible with less compute resources, we employ low-rank adaptation (LoRA) and gradient check-pointing as parameter-efficient strategies. The adapted models are then benchmarked on a subset of the COLE suite of French tasks Beauchemin et al. (2025a), including both Québec French and prestige French benchmarks, to assess their ability to balance dialectal adaptation while retaining general competence in French.
@@
    We demonstrate a compute-efficient CPT pipeline (LoRA + gradient checkpointing) that updates $\leq\!1\%$ of parameters while sustaining throughput on modest hardware (see Section 4). We show that it is possible to adapt an LLM to a regional dialect using only a very small corpus of 86M tokens in our experiments, which is significantly below typical CPT budgets.
@@ 2.1.  Continual Pre-training
Continual Pre-training continues training a model for its original self-supervised objective (e.g., masked or causal language modeling) using a new domain-specific corpus. The objective is to extend the model’s capabilities using additional, unlabeled datasets. CPT differs from supervised fine-tuning, which uses smaller task-specific labeled datasets.
Although CPT is related to continual learning (CL), it is more appropriate for practical specialization rather than lifelong multi-task learning. CL methods such as regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018), replay (Lopez-Paz and Ranzato, 2022; Chaudhry et al., 2019), and gradient orthogonalization (Farajtabar et al., 2020) mitigate catastrophic forgetting across many tasks. On the other hand, CPT typically adapts to a single new distribution without explicit replay or regularizers, retaining original capability through careful optimization and small learning rates (Gururangan et al., 2020b).
While full-parameter fine-tuning is often prohibitive for LLMs due to compute time and cost, low-rank adaptation (LoRA) (Hu et al., 2021) allows updates to only a small fraction of model parameters while preserving pretrained weights. Recent advances combine LoRA with quantization-aware training (e.g., L4Q (Jeon et al., 2025)) to further reduce costs. Other parameter-efficient fine-tuning (PEFT) techniques such as adapters (Houlsby et al., 2019; Pfeiffer et al., 2021), prefix/prompt tuning (Li and Liang, 2021; Lester et al., 2021), bitfit (Ben Zaken et al., 2022), and learned attention/FFN scaling (Liu et al., 2022) reduce memory and computation costs while achieving competitive performances. For resource-constrained continual pre-training, LoRA and related methods can be combined with mixed precision, gradient checkpointing, and quantization to keep adaptation feasible (Dettmers et al., 2023).
@@
Adapting pretrained LLMs to new linguistic varieties can be framed as domain-adaptive pre-training (DAPT) or language-adaptive pre-training (LAPT), where models are exposed to additional unlabeled target-distribution text (Gururangan et al., 2020b; Howard and Ruder, 2018; Chronopoulou et al., 2019). This improves coverage of distribution-specific lexical items, morphology, and style while preserving general knowledge of the base model. Prior work has applied this to new domains (e.g. biomedical or legal)Lee et al. (2019a); Chalkidis et al. (2020) and low-resource languages (Beltagy et al., 2019; Lee et al., 2019b; Koto et al., 2025; Mansha, 2025).
CPT has become a dominant approach for adapting LLMs to low-resource languages and regional dialects Yuan et al. (2024).
Such systems already exist for a variety of languages, trained on datasets as small as 350M tokens; a listing is provided in Table 1. Recent work also shows that mixing English with the target language during CPT can be critical for preserving downstream abilities Yuan et al. (2024).
@@
Basque (Latxa) 4.2B  Agerri et al. (2024)
@@
Vietnamese (VinaLlama) 500B  VietAI (2023)
@@
French LLMs such as CamemBERT and FlauBERT established strong monolingual baselines (Martin et al., 2020; Le et al., 2020), while multilingual encoders (e.g., XLM-RoBERTa) remain competitive for French (Conneau et al., 2020). However, these are trained on the prestige dialect of French used in France. The Québec dialect of French, or Québécois, differs from the prestige dialect in orthography, phonology-to-orthography conventions, anglicisms, idioms, and code-switching patterns. In addition, Québécois resources are comparatively scarce, compared to prestige French or to other languages listed in Table 1. In this case, this makes CPT-based dialectal adaptation valuable.
@@ 3.1.  Data sources
The data sources we used are reported in Table 2. For convenience, we can distinguish two categories: sources that include formal texts (numbers 1 to 3 in Table 2) which make up 60% (52.45M tokens) of our corpus, and those that include informal (spoken or user-generated) texts (numbers 4 to 9) which make up the remaining 40% (34.12M tokens). These sources are described below:
@@
    Public-domain literary texts in Québec French, sourced from the Bibliothèque électronique du Québec (BEQ). Primarily literary text written in normative Québec French. This source offers diachronic coverage of historical Québecois literary language, enriching the corpus with stylistic and temporal diversity.
@@
    Articles linked to the Québec portal of Wikipedia, covering Québec topics and presumed to be primarily written by Québécois people in formal prose. This source contributes broad, topic-specific coverage written in Québec French.
@@
    News articles from the Québec City newspaper, Le Soleil, written in formal journalistic prose. This source offers high-quality formal written text covering a variety of current events.
@@
    Interview transcriptions from the Centre de recherche interuniversitaire sur le français en usage au Québec (CRIFUQ). It is our only example of spontaneous speech in informal Québec French. It thus provides examples of features typical of Québec French speech (e.g., ellipsis, disfluencies, phonology-to-orthography variation).
@@
Overall, our Québec French corpus spans 86.57M tokens, which is significantly below typical CPT budgets used for dialectal/low-resource adaptations, as shown in Table 1.
@@ 4.  Training
To reduce the cost of CPT, we employed LoRA (Hu et al., 2021) instead of updating all model parameters. Following common practice, we targeted the attention projections ($q$, $k$, $v$, $o$) and feed-forward layers (up, down, gate) with rank $r=16$, $\alpha=32$, and dropout of $0.1$. Trainable parameters are cast to float32, while frozen parameters remain in float16. Additionally, we enable gradient checkpointing to further lower memory requirements. This strategy yielded a trainable parameter ratio of around $1\%$ of the full model, making dialectal adaptation feasible on modest hardware.
We performed CPT with a causal language modeling (CLM) objective. The model was exposed to our Québec French corpus (see Section 3) in a single pass without domain interleaving or replay of the previously used prestige French training data. Due to computational constraints, sequences were limited to $1024$ tokens with a stride of $512$, producing overlapping chunks that preserve context for long documents. We trained each model for three and six epochs over the corpus using the AdamW optimizer with a weight decay of $0.01$. The learning rate was set to $1\times 10^{-5}$ with cosine decay and a warm-up ratio of $0.1$. To avoid numerical instability during the training process, we clipped gradients at a norm of $1.0$.
@@
where $b$ is the number of sequences per device (micro-batch), $a$ is the gradient-accumulation steps, and $d$ is the number of devices. With a sequence length of $1024$ tokens, the effective batch size in tokens was therefore $\text{EffBatch}_{\text{tok}}=\text{EffBatch}_{\text{seq}}\times 1024$.
@@ 5.1.  Base Models
  • CroissantLLMChat-v0.1 (1.35B), a bilingual (English/French) open-source model. This serves as a strong dialect-aware baseline that is already exposed to French text.
  • Llama-3.2-1B, a lightweight general-purpose model included as a baseline for multilingual models.
  • Llama-3.1-8B, a high-capacity general-purpose model included as a stronger large-scale baseline for comparison.
@@ 5.2.  Evaluation Benchmarks
We evaluate our models on 8 tasks of the COLE French-language benchmark Beauchemin et al. (2025a). To evaluate both the models’ acquisition of Québec French and their retention of general abilities after CPT, we chose 4 Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE and QFrCoRT) and 4 prestige French tasks (AlloCiné, PAWS-X, Fr-BoolQ, and MMS). These 8 tasks are described below:
@@
    QFrCoLA Beauchemin and Khoury (2025): A grammatical acceptability task. The dataset consists of a total of 25,153 individual sentences classified as grammatical or ungrammatical. The sentences are drawn from a normative Québec French resource, and grouped by phenomena (syntax, morphology, semantics, anglicism), with both in-domain and out-of-domain splits. We used the 7,546 test samples for evaluation.
@@
    MMS Brand24-AI (2023): A sentiment analysis task. This task provides text snippets taken from 79 corpora from various domains, and requires identifying the sentiment in a binary- or trinary-class schema, depending on the corpus. This dataset is also multilingual, but we used the French subset of 63,190 test samples.
@@
None of the text in the Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE, QFrCoRT) were built from, or were included in the Québec French unlabeled corpus used for CPT (see Section 3.1).
@@ 6.1.  Québec French Acquisition
Figure 1 shows the perplexity of the three models before (epoch 0) and after each of the six epochs of CPT. To compute perplexity scores, we used 5% of the COLE test sets during training. Because this is training perplexity (with overlap), it likely underestimates true generalization error; we therefore treat it as a proxy for adaptation dynamics rather than a calibrated held-out metric, and leave non-overlapping held-out perplexity to future work. Nonetheless, the results show that all three models start with high training perplexity, consistent with a distribution gap between their pre-training mix and Québec French, and drop sharply after the first epoch, indicating rapid uptake of dialectal patterns under CPT.
Figure 1: Perplexity during CPT training. Lower perplexity indicates better fit to Québec French. All models evaluated on identical held-out corpus.
@@
We can note as well that there is a second normative Québec French task, QFrBLiMP, on which our models actually improve after CPT. The difference is that this task is a binary comparison, where the model is given two sentences and must tell the correct from the incorrect one. Taken together, these results indicate that CPT is making the models better at understanding both normative and slang Québec French: the models perform better at distinguishing normative from incorrect language, but also more accepting of incorrect popular language.
@@ 7.  Conclusion
In this paper, we demonstrated a case study of low-resource dialectal adaptation of LLMs on Québec French. Our results on 8 tasks from the French benchmark suite COLE show that CPT with LoRA achieves substantial dialectal gains and also improves the model performance on prestige French tasks, but only if the model is large enough to support it. Our work also demonstrated that this specialization can be made by updating approximately 1% of model parameters and with using a very small corpus (85M tokens), but that corpus composition has an important impact. Indeed, our corpus being rich in informal and unedited sources, and lacking question-answering sources made the trained models less proficient at distinguishing normatively correct and incorrect text and at question-answering.

### v3

@@ Abstract
Despite the widespread adoption of Large Language Models (LLMs), their strongest capabilities remain largely confined to a small number of high-resource languages for which there is abundant training data. Recently, continual pre-training (CPT) has emerged as a means to fine-tune these models to low-resource regional dialects. In this paper, we study the use of CPT for dialect learning under tight data and compute budgets. Using low-rank adaptation (LoRA) and compute-efficient continual pre-training, we adapt three LLMs to the Québec French dialect using a very small dataset and benchmark them on the COLE suite. Our experiments demonstrate an improvement on the minority dialect benchmarks with minimal regression on the prestige language benchmarks with around 1% of model parameters updated. Analysis of the results demonstrate that gains are highly contingent on corpus composition. These findings indicate that CPT with parameter-efficient fine-tuning (PEFT) can narrow the dialect gap by providing cost-effective and sustainable language resource creation, expanding high-quality LLM access to minority linguistic communities. To support reproducibility and broaden access, we release the first Québec French LLMs on Hugging Face.
@@ 1.  Introduction
Despite their success, most widely-used LLMs are trained on predominantly English datasets. When other languages are included in multilingual training datasets, they are collected from the high-resource prestige dialect of the language and have little coverage of low-resource regional dialects. As a result, they often struggle with local vocabulary, orthographic variants, idiomatic expressions, and code-switching phenomena that are commonly found in regional dialects. This issue, coined the dialect gap Kantharuban et al. (2023), limits their effectiveness for millions of users of these minority language varieties, thereby perpetuating inequities in access to artificial intelligence technologies.
One possible approach to addressing the dialect gap is to train LLMs using regional dialectal data. However, full model training is prohibitively expensive, while conventional fine-tuning approaches can be inefficient or lead to overfitting on relatively small dialectal datasets. Continual Pre-training (CPT) on unlabeled data offers a practical compromise: by exposing models to large volumes of dialect-specific text, CPT enhances regional linguistic coverage without fully discarding the general knowledge encoded during initial pre-training (Gururangan et al., 2020; Sarkar et al., 2022; Lee et al., 2020).
In this paper, we explore the ability and limitations of CPT to adapt LLMs to a regional dialect using a small amount of unlabeled regional texts. As a case study, we consider specifically the case of the Québec regional dialect of French, also called Québécois. To make adaptation feasible with less compute resources, we employ low-rank adaptation (LoRA) and gradient checkpointing as parameter-efficient strategies. The adapted models are then benchmarked on a subset of the COLE suite of French tasks Beauchemin et al. (2025a), including both Québec French and prestige French benchmarks, to assess their ability to balance dialectal adaptation while retaining general competence in French.
@@
    We demonstrate a compute-efficient CPT pipeline (LoRA + gradient checkpointing) that updates $\leq\!1\%$ of parameters while sustaining performance on modest hardware (see Section 4). We show that it is possible to adapt an LLM to a regional dialect using only a very small corpus of 86M tokens in our experiments, which is significantly below typical CPT data budgets.
@@ 2.1.  Continual Pre-training
Continual Pre-training continues training a model for its original self-supervised objective (e.g., masked or causal language modeling) using an out-of-domain corpus. The objective is to extend the model’s capabilities using additional, unlabeled datasets. CPT differs from supervised fine-tuning, which uses smaller task-specific labeled datasets.
Although CPT is related to continual learning (CL), it is more appropriate for practical specialization rather than lifelong multi-task learning. CL methods such as regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018), replay (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2019), and gradient orthogonalization (Farajtabar et al., 2020) mitigate catastrophic forgetting across many tasks. On the other hand, CPT typically adapts to a single new distribution without explicit replay or regularizers, retaining the model’s original capability through careful optimization and small learning rates (Gururangan et al., 2020).
While full-parameter fine-tuning is often prohibitive for LLMs due to compute time and cost, low-rank adaptation (LoRA) (Hu et al., 2022) allows updates to only a small fraction of model parameters while preserving pretrained weights. Recent advances combine LoRA with quantization-aware training (e.g., L4Q (Jeon et al., 2025)) to further reduce costs. Other parameter-efficient fine-tuning (PEFT) techniques such as adapters (Houlsby et al., 2019; Pfeiffer et al., 2021), prefix/prompt tuning (Li and Liang, 2021; Lester et al., 2021), BitFit (Ben Zaken et al., 2022), and learned attention/FFN scaling (Liu et al., 2022) reduce memory and computation costs while achieving competitive performances. For resource-constrained continual pre-training, LoRA and related methods can be combined with mixed precision, gradient checkpointing, and quantization to keep adaptation feasible (Dettmers et al., 2023).
@@
Adapting pretrained LLMs to new linguistic varieties can be framed as domain-adaptive pre-training (DAPT) or language-adaptive pre-training (LAPT), where models are exposed to additional unlabeled target-distribution text (Gururangan et al., 2020; Howard and Ruder, 2018; Chronopoulou et al., 2019). This improves coverage of distribution-specific lexical items, morphology, and style while preserving general knowledge of the base model. Prior work has applied this to new domains (e.g. biomedical or legal) Lee et al. (2020); Chalkidis et al. (2020) and low-resource languages (Beltagy et al., 2019; Koto et al., 2025; Mansha, 2025).
CPT has become a dominant approach for adapting LLMs to low-resource languages and regional dialects Elhady et al. (2025).
Such systems have been developed for a variety of languages and dialects, trained on datasets as small as 350M tokens. Recent work also shows that mixing English with the target language during CPT can be critical for preserving downstream abilities Elhady et al. (2025).
@@
Basque (Latxa) 4.2B  Etxaniz et al. (2024)
@@
Vietnamese (VinaLlama) 500B  VietAI Research (2023)
@@
French LLMs such as CamemBERT and FlauBERT established strong monolingual baselines (Martin et al., 2020; Le et al., 2020), while multilingual encoders (e.g., XLM-RoBERTa) remain competitive for French (Conneau et al., 2020). However, these are trained on the prestige dialect of French used in France. The Québec dialect of French, or Québécois, differs from the prestige dialect in orthography, phonology-to-orthography conventions, anglicisms, idioms, and code-switching patterns. In addition, Québécois resources are scarce, compared to prestige French or to other languages listed in Table 1. These characteristics make CPT a viable option for dialectal adaptation.
@@ 3.1.  Data sources
The data sources we used are reported in Table 2. We can distinguish two categories: sources that include formal texts (numbers 1 to 3 in Table 2) which make up 60% (52.45M tokens) of our corpus, and those that include informal (spoken or user-generated) texts (numbers 4 to 9) which make up the remaining 40% (34.12M tokens). These sources are described below:
@@
    Public-domain literary texts in Québec French, sourced from the Bibliothèque électronique du Québec (BEQ). These are primarily literary texts written in normative Québec French. This source offers diachronic coverage of historical Québecois literary language, enriching the overall corpus with stylistic and temporal diversity.
@@
    Articles retrieved from the Québec portal of Wikipedia, covering Québec topics and presumed to be primarily written by Québécois people in formal prose. This source contributes broad, topic-specific coverage written in Québec French.
@@
    News articles from the Québec City newspaper, Le Soleil, written in formal journalistic prose. This source offers high-quality formal written texts covering a variety of current events.
@@
    Interview transcriptions from the Centre de recherche interuniversitaire sur le français en usage au Québec (CRIFUQ). It is our only collection of spontaneous speech in informal Québec French. It thus provides examples of features typical of Québec French speech (e.g., ellipsis, disfluencies, phonology-to-orthography variation).
@@
Overall, our Québec French corpus spans 86.57M tokens, which is significantly below typical CPT data budgets used for dialectal/low-resource adaptations, as shown in Table 1.
@@ 4.  Training
To reduce the cost of CPT, we employed LoRA (Hu et al., 2022) instead of updating all model parameters. Following common practice, we targeted the attention projections ($q$, $k$, $v$, $o$) and feed-forward layers (up, down, gate) with rank $r=16$, $\alpha=32$, and dropout of $0.1$. Trainable parameters are cast to float32, while frozen parameters remain in float16. Additionally, we enabled gradient checkpointing to further lower memory requirements. This strategy yielded a trainable parameter ratio of around $1\%$ of the full model, making dialectal adaptation feasible on modest hardware.
We performed CPT with a causal language modeling (CLM) objective. The model was exposed to our Québec French corpus (see Section 3) in a single pass without domain interleaving or replay of the previously used prestige French training data. Due to computational constraints, sequences were limited to $1024$ tokens with a stride of $512$, producing overlapping chunks that preserve context for long documents. We trained each model for 3 and 6 epochs over the corpus using the AdamW optimizer with a weight decay of $0.01$. The learning rate was set to $1\times 10^{-5}$ with cosine decay and a warm-up ratio of $0.1$. To avoid numerical instability during the training process, we clipped gradients at a norm of $1.0$.
@@
where $b$ is the number of sequences per device (micro-batch), $a$ is the gradient-accumulation steps, and $d$ is the number of devices. With a sequence length of $1024$ tokens, the effective batch size in tokens was therefore $\text{EffBatch}_{\text{tok}}=\text{EffBatch}_{\text{seq}}\times 1024$. See Appendix A for more details.
@@ 5.1.  Base Models
  • CroissantLLMChat-v0.1 (1.35B): a bilingual (English/French) open-source model. This serves as a strong dialect-aware baseline that is already exposed to French text.
  • Llama-3.2-1B: a lightweight general-purpose model included as a baseline for multilingual models.
  • Llama-3.1-8B: a high-capacity general-purpose model included as a stronger large-scale multilingual baseline for comparison.
@@ 5.2.  Evaluation Benchmarks
We evaluated our models on 8 tasks of the COLE French-language benchmark Beauchemin et al. (2025a). To evaluate both the models’ acquisition of Québec French and their retention of general abilities after CPT, we chose 4 Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE and QFrCoRT) and 4 prestige French tasks (AlloCiné, PAWS-X, Fr-BoolQ, and MMS). These 8 tasks are described below:
@@
    QFrCoLA Beauchemin and Khoury (2025): A grammatical acceptability task. The dataset consists of a total of 25,153 individual sentences classified as grammatical or ungrammatical. The sentences are drawn from a normative Québec French resource, and grouped by phenomena (syntax, morphology, semantics, anglicism). We used the 7,546 test samples for evaluation.
@@
    MMS Augustyniak et al. (2023): A sentiment analysis task. This task provides text snippets taken from 79 corpora from various domains, and requires identifying the sentiment in a binary- or trinary-class schema, depending on the corpus. This dataset is also multilingual, but we used the French subset of 63,190 test samples.
@@
None of the texts in the Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE, QFrCoRT) were built from, or were included in the Québec French unlabeled corpus used for CPT (see Section 3.1).
@@ 6.1.  Québec French Acquisition
Figure 1 shows the perplexity of the three models before (epoch 0) and after each of the six epochs of CPT. To compute perplexity scores, we used 5% of the COLE test sets during training. Because this is training perplexity (with overlap), it likely underestimates true generalization error. We therefore treat this perplexity as a proxy for adaptation dynamics rather than a calibrated held-out metric, and leave non-overlapping held-out perplexity to future work. Nonetheless, the results show that all three models start with high training perplexity, consistent with a distribution gap between their pre-training mix and Québec French, and drop sharply after the first epoch, indicating rapid uptake of dialectal patterns under CPT.
Figure 1: Perplexity during CPT training. Lower perplexity indicates better fit to Québec French. All models are evaluated on identical held-out corpus.
@@
We can note as well that there is a second normative Québec French task, QFrBLiMP, on which our models actually improve after CPT. The difference is that this task is a binary comparison, where the model is given two sentences and must identify the correct from the incorrect one. Taken together, these results indicate that CPT is making the models better at understanding both normative and slang Québec French: the models perform better at distinguishing normative from incorrect language, but also more accepting of incorrect popular language.
@@ 7.  Conclusion
In this paper, we demonstrated a case study of low-resource dialectal adaptation of LLMs on Québec French. Our results on 8 tasks from the French benchmark suite COLE show that CPT with LoRA achieves substantial dialectal gains and also improves the model performance on prestige French tasks, but only if the model is large enough to support it. Our work also demonstrated that this specialization can be made by updating $\leq\!1\%$ of model parameters and with using a very small corpus (85M tokens), but that corpus composition has an important impact. Indeed, our corpus being rich in informal and unedited sources, and lacking question-answering sources made the trained models less proficient at distinguishing normatively correct and incorrect text and at question-answering.